# Decentralized projected Riemannian gradient methodKangkang Deng, Jiang Hu

### v1

@@ Abstract
This work addresses the problem of decentralized optimization on a compact submanifold within a communication network comprising $n$ nodes. Each node is associated with a smooth, non-convex local cost function, and the collective objective is to minimize the sum of these local cost functions. We focus on an online scenario where local data arrives continuously in a streaming fashion, eliminating the necessity for complete data storage. To tackle this problem, we introduce a novel algorithm, the Decentralized Projected Riemannian Stochastic Recursive Momentum (DPRSRM) method. Our approach leverages hybrid local stochastic gradient estimators and utilizes network communication to maintain a consensus on the global gradient. Notably, DPRSRM attains an oracle complexity of $\mathcal{O}(\epsilon^{-\frac{3}{2}})$, which surpasses the performance of existing methods with complexities no better than $\mathcal{O}(\epsilon^{-2})$. Each node in the network requires only $\mathcal{O}(1)$ gradient evaluations per iteration, avoiding the need for large batch gradient calculations or restarting procedures. Finally, we validate the superior performance of our proposed algorithm through numerical experiments, including applications in principal component analysis and low-rank matrix completion, demonstrating its advantages over state-of-the-art approaches.
@@ 1 Introduction
where $n$ is the number of nodes, $f_{i}$ is the smooth nonconvex local objective at the $i$-th node, and $\mathcal{M}$ is a (nonconvex) compact smooth submanifold embedded in $\mathbb{R}^{d\times r}$ with the extrinsic dimensions $(d,r)$, e.g., the Stiefel manifold ${\rm St}(d,r):=\{x\in\mathbb{R}^{d\times r}:x^{\top}x=I_{r}\}$. Problem (1.1) is prevalent in machine learning, signal processing, and deep learning, see, e.g., the principal component analysis [35], the low-rank matrix completion [4, 20, 12], the low-dimensional subspace learning [1, 22], and the deep neural networks with batch normalization [8, 17]. One challenge in solving (1.1) comes from the nonconvexity of the manifold constraint, which causes difficulty in achieving the consensus [7, 12, 19].
@@ 1.1 Related works
Several studies have focused on the finite-sum setting of problem (1.1), where $f_{i}=\frac{1}{m}\sum_{r=1}^{m_{i}}f_{i,r}$. [7] propose a decentralized Riemannian stochastic gradient descent method. By combining the variable sample size gradient approximation method with the gradient tracking dynamic, [38] propose a distributed Riemannian stochastic optimization algorithm on the Stiefel manifold. Although both methods can also be used in the online setting, the oracle complexities are $\mathcal{O}(\epsilon^{-2})$, which is not optimal. It is worth noting that the decentralized variance reduced method [33] has been studied. However, they need to periodically calculate the full gradient, making it not suitable for the online setting.
@@ 2.1 Stationary point
$\frac{1}{n}\sum_{i=1}^{n}\|x_{i}-\bar{x}\|^{2}\leq\epsilon\quad{\rm and}\quad\|\mathrm{grad}f(\bar{x})\|^{2}\leq\epsilon.$
@@ 2.2 Proximal smoothness
Proximal smoothness is an effective tool for addressing the nonconvex nature of manifold constraints within decentralized optimization settings [12]. Define the distance from a point $x\in\mathbb{R}^{d\times r}$ to the manifold $\mathcal{M}$ and the nearest-point projection of $x$ onto $\mathcal{M}$ as $\text{dist}(x,\mathcal{M}):=\inf_{y\in\mathcal{M}}\|y-x\|\;\;\text{and}\;\;\mathcal{P}_{\mathcal{M}}(x):=\arg\min_{y\in\mathcal{M}}\|y-x\|,$ respectively. For a given number $R>0$, the $R$-tube around $\mathcal{M}$ is defined as the set $U_{\mathcal{M}}(R):=\{x:\text{dist}(x,\mathcal{M})<R\}.$ A closed set $\mathcal{M}$ is said to be $R$-proximally smooth if the projection $\mathcal{P}_{\mathcal{M}}(x)$ is unique whenever $\text{dist}(x,\mathcal{M})<R$. Following [9], an $R$-proximally smooth set $\mathcal{M}$ satisfies that for any real ${\delta}\in(0,R)$,
@@
where $\bar{U}_{\mathcal{M}}(\delta):=\{x:\text{dist}(x,\mathcal{M})\leq\delta\}$. For instance, the Stiefel manifold is known to be a 1-proximally smooth set [2].
@@ 3.1 Problem setup
Let us start with the following assumptions on problem (1.1). For the manifold $\mathcal{M}$, it is shown that any compact $C^{2}$-submanifolds in Euclidean space belong to proximally smooth set [2, 9, 11]. Without loss of generality, we assume that $\mathcal{M}$ is $R$-proximally smooth. For the objective function $f_{i}$, we make the following assumption.
@@
Each objective function $f_{i}$ is gradient Lipschitz continuous with modulus $L_{f}$ on the convex hull of $\mathcal{M}$, denoted by ${\rm conv}(\mathcal{M})$. Moreover, the objective function $f(\mathbf{x})$ has an optimal value $f_{*}$ over $\mathcal{M}^{n}$.
@@
(3.1) $f_{i}(y)-f_{i}(x)\leq\left<\mathrm{grad}f_{i}(x),y-x\right>+\frac{L_{g}}{2}\|y-x\|^{2},~{}i\in[n],$
@@
To measure the oracle complexity of our algorithm, we give the definition of a stochastic first-order oracle (SFO) for (1.1).
@@ 3.2 The Algorithm
In this subsection, we introduce the Decentralized Projected Riemannian Stochastic Recursive Momentum (DPRSRM) method for addressing (1.1), along with its associated convergence results. Inspired by the notable effectiveness of variance reduction and gradient tracking in decentralized frameworks [33, 12, 7], we seek a novel combination of variance reduction and gradient tracking for decentralized online problems on compact submanifolds for improving the oracle complexity. In particular, we focus on the following stochastic gradient estimator using the momentum technique introduced in [10, 15]:
@@
A crucial advantage of gradient tracking-type methods lies in the applicability of the use of a constant step size $\alpha$. Since $s_{i,k}$ may not remain in in the tangent space $T_{x_{i,k}}\mathcal{M}$, we introduce the projected variable $v_{i,k}$ and update the new iterate $x_{i,k+1}$ as follows:
(3.10) $x_{i,k+1}=\mathcal{P}_{\mathcal{M}}(\sum_{j=1}^{n}W_{ij}x_{j,k}-\alpha v_{i,k}),~{}i\in[n].$
For ease of analysis, we stack variables in each node and rewrite (3.6), (3.9) and (3.10) as
@@
where $\mathbf{\xi}_{k}=\{\xi_{i,k}\}_{i\in\mathcal{V}}$ and $\mathcal{P}_{T_{\mathbf{x}}\mathcal{M}^{n}}:=\mathcal{P}_{T_{x_{1,k}}\mathcal{M}}\times\dots\times\mathcal{P}_{T_{x_{n,k}}\mathcal{M}}$. The overall algorithm is given in Algorithm 1.
@@
0:  Initial point $\mathbf{x}_{0}=\bar{x}_{0}\in\mathcal{M}$, $\mathrm{\bf s}_{-1}=\mathbf{d}_{-1}=\mathbf{0}$, $\alpha,\tau$.
@@ 3.3 Main results
Let $\{\mathbf{x}_{k}\}_{k}$ be the sequence generated by Algorithm 1 with $B\geq L$. Suppose that Assumptions 3.1-3.4 hold. Let $\alpha=\frac{1}{K}$, $\tau=\frac{1}{K^{2/3}}$ and $\delta\leq\min\{\frac{R}{2},\frac{R(1-\sigma_{2})}{2}\}$. If $\mathbf{x}_{0}\in\mathcal{N}(\delta)$, and $K\geq\max\left\{4L,\sqrt{nD}/\delta,\frac{R\sqrt{nD}}{(R(1-\sigma_{2})-2\delta)\delta}\right\}$, it holds that
(3.17) $\frac{1}{n}\|\bar{\mathbf{x}}_{k}-\mathbf{x}_{k}\|^{2}$ $\leq\frac{C}{K^{2/3}},$
and
(3.18) $\min_{0\leq k\leq K}\|\mathrm{grad}f(\bar{x}_{k})\|^{2}\leq\frac{\Gamma_{1}}{K^{2/3}}+\frac{\Gamma_{2}}{K}+\frac{6\rho_{2}\nu^{2}}{K^{4/3}},$
@@
(3.19) $\mathcal{K}_{1}:$ $=(C+\Gamma_{1})^{1.5}\epsilon^{-\frac{3}{2}},\mathcal{K}_{2}:=\Gamma_{2}\epsilon^{-1},\mathcal{K}_{3}:=(6\rho_{2}\nu^{2})^{3/4}\epsilon^{-\frac{3}{4}}.$
@@ 4.1 Consensus error
Suppose that Assumption 3.3 holds. Let $x_{i,k}$ be generated by Algorithm 1. Let $\alpha$ and $\delta$ satisfy (3.13). If $\mathbf{x}_{0}\in\mathcal{N}(\delta)$, then for any $k\geq 1$, it holds that $\mathbf{x}_{k}\in\mathcal{N}(\delta)$ and
@@ 4.2 Optimality error
We also have the following gradient estimation error bound.
@@ 5 Numerical experiments
In this section, we compare our proposed DPRSRM with DRSGD in [7] and DRPGD in [12] on decentralized principal component analysis and low-rank matrix completion problem. It is important to note that the original DRPGD is a deterministic algorithm that utilizes the local full gradient. To adapt it to the online setting, we replace the full gradient with a stochastic gradient for the local updates. The numerical results on decentralized low-rank matrix completion are provided in the supplementary material.
@@ 5.1.1 Synthetic dataset
Firstly, We test all algorithms with different network graphs, namely, Ring, ER $p=0.3$, and ER $p=0.6$. The results are shown in Figure 1. It can be seen that there is not much difference among different graphs except for the consensus error. This is consistent with the existing results for DRSGD [7] and DPRGD [12]. Secondly, Figure 2 presents a comparison among the three algorithms. Our DPRSRM outperforms the other two, with DRSGD and DPRGD showing comparable performance.
@@ 5.1.2 Mnist dataset
To evaluate the efficiency of the proposed DPRSRM, we also conduct numerical tests on the Mnist dataset [21]. The training images consist of 60000 handwritten images of size $32\times 32$ and are used to generate $A_{i}$’s. We first normalize the data matrix by dividing 255 and randomly split the data into $n=8$ nodes with equal cardinality. Then, each node holds a local matrix $A_{i}$ of dimension $\frac{60000}{n}\times 784$. We compute the first 5 principal components, i.e., $d=784,r=5$.
@@ 6 Conclusions and Limitations
This develop a decentralized projection Riemannian stochastic recursive momentum method by assuming that each node has access to a stochastic first-order oracle. Our algorithm leverages local hybrid variance reduction and gradient tracking to achieve a lower oracle complexity compared with the existing online methods. It requires only $\mathcal{O}(1)$ gradient evaluations per iteration for each local node and does not require restarting with a large batch gradient.
Limitations. Our paper motivates several important questions for future research. The lack of closed-form solutions for the projection operator on certain manifolds suggests a need to explore methods for approximate projection calculation. Additionally, our step-size selection depends on the proximal smoothness constant $R$, emphasizing the importance of accurately estimating $R$ for specific manifolds.
@@ Appendix B Proof of Section 4
$a_{3}\leq$ $\frac{4(8Q+\sqrt{n}L_{2}+M_{2})^{2}}{n^{2}}\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|^{4}+\frac{16Q^{2}\alpha^{4}}{n^{2}}\|{\bf v}_{k}\|^{4}+4\alpha^{2}\|\hat{v}_{k}\|^{2}$
$+\frac{4M_{2}^{2}}{n^{2}}\|\mathbf{x}_{k+1}-\bar{\mathbf{x}}_{k+1}\|^{4}$
$\leq$ $\frac{4(8Q+\sqrt{n}L_{2}+M_{2})^{2}}{n^{2}}\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|^{4}+\frac{16Q^{2}\alpha^{4}}{n^{2}}\|{\bf s}_{k}\|^{4}+\frac{4\alpha^{2}}{n}\|\mathrm{\bf s}_{k}\|^{2}$
$+\frac{4M_{2}^{2}}{n^{2}}\|\mathbf{x}_{k+1}-\bar{\mathbf{x}}_{k+1}\|^{4}.$
@@
(B.9) $a_{4}=$ $\mathrm{grad}f(\bar{x}_{k}),\frac{1}{n}\sum_{i=1}^{n}\alpha(s_{i,k}-v_{i,k})\right>$
$=\frac{\alpha}{n}\sum_{i=1}^{n}\left<\mathrm{grad}f(\bar{x}_{k})-\mathcal{P}_{T_{x_{i,k}}\mathcal{M}}(\mathrm{grad}f(\bar{x}_{k})),s_{i,k}-v_{i,k}\right>$
@@
(B.14) $\mathbb{E}[\|\mathbf{d}_{k+1}-\mathbf{d}_{k}\|^{2}]$
$\leq 3\tau^{2}\mathbb{E}[\|\mathrm{grad}f(\mathbf{x}_{k})-\mathbf{d}_{k}\|^{2}]+24L^{2}\mathbb{E}[\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|^{2}]+12L^{2}\alpha^{2}\mathbb{E}[\|\mathbf{v}_{k}\|^{2}]+3n\tau^{2}\nu^{2}.$
@@
(B.15) $\mathbb{E}[\|{\bf s}_{k+1}-\hat{\bf d}_{k+1}\|^{2}]\leq(1+\zeta)\sigma_{2}^{2}\mathbb{E}[\|{\bf s}_{k}-\hat{\bf d}_{k}\|^{2}]+(1+\frac{1}{\zeta})\mathbb{E}[\|\mathbf{d}_{k+1}-\mathbf{d}_{k}\|^{2}]\leq(1+\zeta)\sigma_{2}^{2}\mathbb{E}[\|{\bf s}_{k}-\hat{\bf d}_{k}\|^{2}]$
$+(1+\frac{1}{\zeta})(3\tau^{2}\mathbb{E}[\|\mathrm{grad}f(\mathbf{x}_{k})-\mathbf{d}_{k}\|^{2}]+24L^{2}\mathbb{E}[\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|^{2}]+12L^{2}\alpha^{2}\mathbb{E}[\|\mathbf{v}_{k}\|^{2}]+3n\tau^{2}\nu^{2})$
@@
(B.16) $\mathbb{E}[\|{\bf s}_{k+1}-\hat{\bf d}_{k+1}\|^{2}]\leq\frac{1+\sigma_{2}^{2}}{2}\mathbb{E}[\|{\bf s}_{k}-\hat{\bf d}_{k}\|^{2}]$
$+3\frac{1+\sigma_{2}^{2}}{1-\sigma_{2}^{2}}(\tau^{2}\mathbb{E}[\|\mathrm{grad}f(\mathbf{x}_{k})-\mathbf{d}_{k}\|^{2}]+8L^{2}\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|^{2}+4L^{2}\alpha^{2}\|\mathbf{v}_{k}\|^{2}+n\tau^{2}\nu^{2}).$

### v2

@@ Abstract
This paper studies decentralized optimization over a compact submanifold within a communication network of $n$ nodes, where each node possesses a smooth non-convex local cost function, and the goal is to jointly minimize the sum of these local costs. We focus particularly on the online setting, where local data is processed in real-time as it streams in, without the need for full data storage. We propose a decentralized projected Riemannian stochastic recursive momentum (DPRSRM) method that employs local hybrid stochastic gradient estimators and uses the network to track the global gradient. DPRSRM achieves an oracle complexity of $\mathcal{O}(\epsilon^{-\frac{3}{2}})$, outperforming existing methods that have at most $\mathcal{O}(\epsilon^{-2})$ complexity. Our method requires only $\mathcal{O}(1)$ gradient evaluations per iteration for each local node and does not require restarting with a large batch gradient. Furthermore, we demonstrate the effectiveness of our proposed methods compared to state-of-the-art ones through numerical experiments on principal component analysis problems and low-rank matrix completion.
@@ 1 Introduction
where $n$ is the number of nodes, $f_{i}$ is the smooth nonconvex local objective at the $i$-th node, and $\mathcal{M}$ is a (nonconvex) compact smooth submanifold embedded in $\mathbb{R}^{d\times r}$ with the extrinsic dimensions $(d,r)$, e.g., the Stiefel manifold ${\rm St}(d,r):=\{x\in\mathbb{R}^{d\times r}:x^{\top}x=I_{r}\}$. Problem (1.1) is prevalent in machine learning, signal processing, and deep learning, see, e.g., the principal component analysis [37], the low-rank matrix completion [4, 22, 12], the low-dimensional subspace learning [1, 24], and the deep neural networks with batch normalization [8, 19]. One challenge in solving (1.1) comes from the nonconvexity of the manifold constraint, which causes difficulty in achieving the consensus [7, 12, 21, 13].
@@ 1.1 Related works
Several studies have focused on the finite-sum setting of problem (1.1), where $f_{i}=\frac{1}{m}\sum_{r=1}^{m_{i}}f_{i,r}$. [7] propose a decentralized Riemannian stochastic gradient descent method. By combining the variable sample size gradient approximation method with the gradient tracking dynamic, [40] propose a distributed Riemannian stochastic optimization algorithm on the Stiefel manifold. Although both methods can also be used in the online setting, the oracle complexities are $\mathcal{O}(\epsilon^{-2})$, which is not optimal. It is worth noting that the decentralized variance reduced method [35] has been studied. However, they need to periodically calculate the full gradient, which is not suitable for the online setting.
@@ 2.1 Stationary point
$\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}\|x_{i}-\bar{x}\|^{2}\right]\leq\epsilon\quad{\rm and}\quad\mathbb{E}[\|\mathrm{grad}f(\bar{x})\|^{2}]\leq\epsilon.$
@@ 2.2 Proximal smoothness
Proximal smoothness is an effective tool for addressing the nonconvex nature of manifold constraints within decentralized optimization settings [12]. Define the distance from a point $x\in\mathbb{R}^{d\times r}$ to the manifold $\mathcal{M}$ and the nearest-point projection of $x$ onto $\mathcal{M}$ as $\text{dist}(x,\mathcal{M}):=\inf_{y\in\mathcal{M}}\|y-x\|\;\;\text{and}\;\;\mathcal{P}_{\mathcal{M}}(x):=\arg\min_{y\in\mathcal{M}}\|y-x\|,$ respectively. For a given number $R>0$, the $R$-tube around $\mathcal{M}$ is defined as the set $U_{\mathcal{M}}(R):=\{x:\text{dist}(x,\mathcal{M})<R\}.$
A closed set $\mathcal{M}$ is said to be $R$-proximally smooth if the projection $\mathcal{P}_{\mathcal{M}}(x)$ is unique whenever $\text{dist}(x,\mathcal{M})<R$. Following [9], an $R$-proximally smooth set $\mathcal{M}$ satisfies that for any real ${\delta}\in(0,R)$,
@@
where $\bar{U}_{\mathcal{M}}(\delta):=\{x:\text{dist}(x,\mathcal{M})\leq\delta\}$. For instance, the Stiefel manifold is known to be a 1-proximally smooth set [2], which also provides additional examples of rank manifolds along with their specific proximal smoothness radii.
@@ 3.1 Problem setup
Let us start with assumptions on problem (1.1), based on the fact that any compact $C^{2}$-submanifolds in Euclidean space belong to proximally smooth set [2, 9, 11].
@@
Assume that $\mathcal{M}$ is proximally smooth with radius $R$. Each objective function $f_{i}$ is gradient Lipschitz continuous with modulus $L_{f}$ on the convex hull of $\mathcal{M}$, denoted by ${\rm conv}(\mathcal{M})$. Moreover, the objective function $f(\mathbf{x})$ has an optimal value $f_{*}$ over $\mathcal{M}^{n}$.
@@
(3.1) $f_{i}(y)-f_{i}(x)\leq\left<\mathrm{grad}f_{i}(x),y-x\right>+\frac{L_{g}}{2}\|y-x\|^{2},~{}i\in[n].$
@@
To measure the oracle complexity, we give the definition of a stochastic first-order oracle (SFO) for (1.1).
@@ 3.2 The Algorithm
In this subsection, we introduce the Decentralized Projected Riemannian Stochastic Recursive Momentum (DPRSRM) method for addressing (1.1), along with its associated convergence results. Inspired by the notable effectiveness of variance reduction and gradient tracking in decentralized frameworks [35, 12, 7], we seek a novel combination of variance reduction and gradient tracking for decentralized online problems on compact submanifolds for improving the oracle complexity. In particular, we focus on the following stochastic gradient estimator using the momentum technique introduced in [10, 17, 14]:
@@
A crucial advantage of gradient tracking-type methods lies in the applicability of the use of a constant step size $\alpha$. Since $s_{i,k}$ may not remain in the tangent space $T_{x_{i,k}}\mathcal{M}$, we introduce its tangent space projection $v_{i,k}=\mathcal{P}_{T_{x_{i,k}}(s_{i,k})}$ and update the new iterate $x_{i,k+1}$ as follows:
(3.10) $x_{i,k+1}=\mathcal{P}_{\mathcal{M}}(\sum_{j=1}^{n}W_{ij}x_{j,k}-\alpha v_{i,k}),~{}i\in[n],$
where the projection-based average [12, 20], i.e., $\mathcal{P}_{\mathcal{M}}(\sum_{j=1}^{n}W_{ij}x_{j,k})$, is adopted to ensure the decrease in the consensus error among nodes. For ease of analysis, we stack variables in each node and rewrite (3.6), (3.9) and (3.10) as
@@
where $\mathcal{P}_{T_{\mathbf{x}}\mathcal{M}^{n}}:=\mathcal{P}_{T_{x_{1,k}}\mathcal{M}}\times\dots\times\mathcal{P}_{T_{x_{n,k}}\mathcal{M}}$ and $\mathbf{\xi}_{k}=\{\xi_{i,k}\}_{i\in\mathcal{V}}$. The overall algorithm is given in Algorithm 1.
@@
0:  Initial point $\mathbf{x}_{0}=\bar{\mathbf{x}}_{0}\in\mathcal{M}$, $\mathrm{\bf s}_{-1}=\mathbf{d}_{-1}=\mathbf{0}$, $\alpha,\tau$.
@@ 3.3 Main results
Let $\{\mathbf{x}_{k}\}_{k}$ be the sequence generated by Algorithm 1 with $B\geq L$. Suppose that Assumptions 3.1-3.4 hold. Let $\alpha=\frac{1}{K^{1/3}}$, $\tau=\frac{1}{K^{2/3}}$ and $\delta\leq\min\{\frac{R}{2},\frac{R(1-\sigma_{2})}{2}\}$. If $\mathbf{x}_{0}\in\mathcal{N}(\delta)$, and $K\geq\max\left\{4L,\sqrt{nD}/\delta,\frac{R\sqrt{nD}}{(R(1-\sigma_{2})-2\delta)\delta}\right\}$, it holds that
(3.17) $\mathbb{E}\left[\frac{1}{n}\|\bar{\mathbf{x}}_{K}-\mathbf{x}_{K}\|^{2}\right]$ $\leq\frac{C}{K^{2/3}},$
$\min_{0\leq k\leq K}\mathbb{E}[\|\mathrm{grad}f(\bar{x}_{k})\|^{2}]$ $\leq\frac{\Gamma_{1}}{K^{2/3}}+\frac{\Gamma_{2}}{K}+\frac{6\rho_{2}\nu^{2}}{K^{4/3}},$
@@
(3.18) $\mathcal{K}_{1}:$ $=(C+\Gamma_{1})^{1.5}\epsilon^{-\frac{3}{2}},\mathcal{K}_{2}:=\Gamma_{2}\epsilon^{-1},$
$\mathcal{K}_{3}:$ $=(6\rho_{2}\nu^{2})^{3/4}\epsilon^{-\frac{3}{4}}.$
@@ 4.1 Consensus error
Suppose that Assumptions 3.1 and 3.3 hold. Let $x_{i,k}$ be generated by Algorithm 1. Let $\alpha$ and $\delta$ satisfy (3.13). If $\mathbf{x}_{0}\in\mathcal{N}(\delta)$, then for any $k\geq 1$, it holds that $\mathbf{x}_{k}\in\mathcal{N}(\delta)$ and
@@ 4.2 Optimality error
We also have the gradient estimation error bound.
@@ 5 Numerical experiments
In this section, we compare our proposed DPRSRM with DRSGD in [7] and DRPGD in [12] on decentralized principal component analysis. It is important to note that the original DRPGD is a deterministic algorithm that utilizes the local full gradient. To adapt it to the online setting, we replace the full gradient with a stochastic gradient for the local updates. The numerical results on decentralized low-rank matrix completion are provided in the supplementary material.
@@ 5.1.1 Synthetic dataset
Firstly, we test all algorithms with different network graphs, namely, Ring, ER $p=0.3$, and ER $p=0.6$. The results are shown in Figure 1. We see that there is not much difference among different graphs except for the consensus error. This is consistent with the existing results for DRSGD [7] and DPRGD [12]. Secondly, Figure 2 presents a comparison among the three algorithms. Our DPRSRM outperforms the other two, with DRSGD and DPRGD showing comparable performance.
@@ 5.1.2 Mnist dataset
We also conduct numerical tests on the Mnist dataset [23]. The training images consist of 60000 handwritten images of size $32\times 32$ and are used to generate $A_{i}$’s. We first normalize the data matrix by dividing 255 and randomly split the data into $n=8$ nodes with equal cardinality. Then, each node holds a local matrix $A_{i}$ of dimension $\frac{60000}{n}\times 784$. We compute the first 5 principal components, i.e., $d=784,r=5$.
@@ 6 Conclusions and Limitations
This work develops a decentralized projection Riemannian stochastic recursive momentum method by assuming that each node has access to a stochastic first-order oracle. Our algorithm leverages local hybrid variance reduction and gradient tracking to achieve a lower oracle complexity compared with the existing online methods. It requires only $\mathcal{O}(1)$ gradient evaluations per iteration for each local node and does not require restarting with a large batch gradient.
Acknowledgements
This effort was supported by the SciAI Center, and funded by the Office of Naval Research (ONR), under Grant Number N00014-23- 1-2729 (J.H.), as well as by the National Natural Science Foundation of China (NSFC) grants 12401419 (K.D.).
@@ Appendix B Proof of Section 4
$a_{3}\leq$ $\frac{4(8Q+\sqrt{n}L_{2}+M_{2})^{2}}{n^{2}}\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|^{4}+\frac{16Q^{2}\alpha^{4}}{n^{2}}\|{\bf v}_{k}\|^{4}+4\alpha^{2}\|\hat{v}_{k}\|^{2}+\frac{4M_{2}^{2}}{n^{2}}\|\mathbf{x}_{k+1}-\bar{\mathbf{x}}_{k+1}\|^{4}$
$\leq$ $\frac{4(8Q+\sqrt{n}L_{2}+M_{2})^{2}}{n^{2}}\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|^{4}+\frac{16Q^{2}\alpha^{4}}{n^{2}}\|{\bf s}_{k}\|^{4}+\frac{4\alpha^{2}}{n}\|\mathrm{\bf s}_{k}\|^{2}+\frac{4M_{2}^{2}}{n^{2}}\|\mathbf{x}_{k+1}-\bar{\mathbf{x}}_{k+1}\|^{4}.$
@@
(B.9) $a_{4}=$ $\mathrm{grad}f(\bar{x}_{k}),\frac{1}{n}\sum_{i=1}^{n}\alpha(s_{i,k}-v_{i,k})\right>=\frac{\alpha}{n}\sum_{i=1}^{n}\left<\mathrm{grad}f(\bar{x}_{k})-\mathcal{P}_{T_{x_{i,k}}\mathcal{M}}(\mathrm{grad}f(\bar{x}_{k})),s_{i,k}-v_{i,k}\right>$
@@
(B.14) $\mathbb{E}[\|\mathbf{d}_{k+1}-\mathbf{d}_{k}\|^{2}]\leq 3\tau^{2}\mathbb{E}[\|\mathrm{grad}f(\mathbf{x}_{k})-\mathbf{d}_{k}\|^{2}]+24L^{2}\mathbb{E}[\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|^{2}]+12L^{2}\alpha^{2}\mathbb{E}[\|\mathbf{v}_{k}\|^{2}]+3n\tau^{2}\nu^{2}.$
@@
(B.15) $\mathbb{E}[\|{\bf s}_{k+1}-\hat{\bf d}_{k+1}\|^{2}]\leq(1+\zeta)\sigma_{2}^{2}\mathbb{E}[\|{\bf s}_{k}-\hat{\bf d}_{k}\|^{2}]+(1+\frac{1}{\zeta})\mathbb{E}[\|\mathbf{d}_{k+1}-\mathbf{d}_{k}\|^{2}]$
$\leq$ $(1+\zeta)\sigma_{2}^{2}\mathbb{E}[\|{\bf s}_{k}-\hat{\bf d}_{k}\|^{2}]+(1+\frac{1}{\zeta})(3\tau^{2}\mathbb{E}[\|\mathrm{grad}f(\mathbf{x}_{k})-\mathbf{d}_{k}\|^{2}]+24L^{2}\mathbb{E}[\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|^{2}]+12L^{2}\alpha^{2}\mathbb{E}[\|\mathbf{v}_{k}\|^{2}]+3n\tau^{2}\nu^{2})$
@@
(B.16) $\mathbb{E}[\|{\bf s}_{k+1}-\hat{\bf d}_{k+1}\|^{2}]$
$\leq$ $\frac{1+\sigma_{2}^{2}}{2}\mathbb{E}[\|{\bf s}_{k}-\hat{\bf d}_{k}\|^{2}]+3\frac{1+\sigma_{2}^{2}}{1-\sigma_{2}^{2}}(\tau^{2}\mathbb{E}[\|\mathrm{grad}f(\mathbf{x}_{k})-\mathbf{d}_{k}\|^{2}]+8L^{2}\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|^{2}+4L^{2}\alpha^{2}\|\mathbf{v}_{k}\|^{2}+n\tau^{2}\nu^{2}).$